OpenStudy (anonymous):

reduce the ratio to lowest terms. a/b ^{2} x^2 to 1^2/b x

OpenStudy (anonymous):



OpenStudy (phi):

you can write the ratio \[ \frac{a}{b^2}x^2 : \frac{a^2}{b} x \] as \[ \frac{ \frac{a}{b^2}x^2 }{ \frac{a^2}{b} x }\] you can change this division to multiplying by the inverse of the bottom fraction. In other words, "flip" the bottom fraction, and multiply: \[ \frac{a}{b^2}x^2 \cdot \frac{b}{a^2 x} \] or \[ \frac{abx^2}{a^2b^2x} \] you can "cancel" like terms from the top and bottom Can you finish ?

OpenStudy (anonymous):

x/ab ??

OpenStudy (phi):

yes, assuming I typed in the correct question.

OpenStudy (anonymous):

yes,, ok tyvm!!!

OpenStudy (phi):

yw